satellite73 (satellite73):

as per estudier On average how many times do you have to toss a coin to get a run of an odd number of heads followed by a tail?

OpenStudy (anonymous):

dang probability

OpenStudy (anonymous):

i thought this was simple when i saw it at first, but now i am thinking it is not

OpenStudy (anonymous):

To be honest, I don't even understand the question

OpenStudy (anonymous):

odd number of head followed by a tail. i take it to mean this. h t or t h t or t t h t h h h t or t h h h t h h t h t

OpenStudy (anonymous):

each as odd number of heads followed by a tail, and i think those are the only possibilities for success in the number of tosses for each case

OpenStudy (anonymous):

2 tosses: h t 3 tosses: t h t 4 tosses: t t h t or h h h t 5 tosses: t h h h t or h h t h t

OpenStudy (anonymous):

then you would multiply and add \[2\times \frac{1}{4}+3\times \frac{1}{8}+ 4\times \frac{1}{8}+5\times \frac{1}{16}+...\] of course all the work is in the ... part

OpenStudy (anonymous):

U have 1/2 squared + fourth + sixth + = 1/3 which is probability to succeed right off the bat so fail means tail after even heads. Then u have to do that 3 times, it is true, but that is trials not tosses. Each trial ends with a tail so you can say that the average number of tosses per trial is 1/half or 2. So 2*3 = 6.

OpenStudy (anonymous):

that is what i thought at first too. but while i was out i was thinking about it and i changed my mind

OpenStudy (anonymous):

Else x = 1/2 (1+x) +1/4(2+x) + 1/4 *2 -> x=6.

OpenStudy (anonymous):

i was thinking you were right. i said 3 and forgot that it was for two tosses and you said six and i thought you were correct, but then i started thinking that we are both wrong. look at what i wrote above for cases of success

OpenStudy (anonymous):

A (small) set of trials would seem to confirm that is nearer to 6 than to 3. Simplest would be to plug it in to a simulator and see..

OpenStudy (anonymous):

ok but i am still looking for a reason because look at the possibilities i wrote above.

OpenStudy (anonymous):

5 tosses: t h h h t or h h t h t the second one is tail after even heads

OpenStudy (anonymous):

yes but it says odd number of head followed by a tail. so h h t h t works for 5 tosses because you have one head followed by one tail. the last two. in other words i think these are the only two possibilities for success on 5 tosses. and that is why i think this is rather complicated

OpenStudy (anonymous):

Then u go on to 5h's and T

OpenStudy (anonymous):

i count two ways for 5 tosses. not one way. how about 6 tosses?

OpenStudy (anonymous):

That's what I just said, 5 h's and t

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

the way i read it, it does not say odd number of heads in a row and and then tails. i just says how long do you have to wait to get a run of and odd number of heads and then a tail

OpenStudy (anonymous):

so for example h h h h h t is one way but so is t t h h h t 6 tosses and the last 4 give you a string of three heads followed by a tail

OpenStudy (anonymous):

also h h t t h t works and so does t t t t h t

OpenStudy (anonymous):

doesn't say what happens on the first bunch of tosses, just that you end up with odd number of heads followed by a tail. so both of those i would count as having a string of odd number of heads followed by a tail

OpenStudy (anonymous):

OK, we keep talking about different things, I am talking about probability of success and u are trying to establish the event space....

OpenStudy (anonymous):

yes but in order to find the probability of success i need to know what a success is. how many ways to do it on two tosses? we both get one way so probability is 1/4

OpenStudy (anonymous):

on three tosses we both get one way t h t so probability is 1/8

OpenStudy (anonymous):

The 1/3 is based on that u start tossing odd heads followed by a tail....

OpenStudy (anonymous):

So that failure is then based on tail after even heads...

OpenStudy (anonymous):

in four tosses i count t t h t h h h t so probability is 1/8 if you don't count t t h t then we will get different answers

OpenStudy (anonymous):

you see the difference yes?

OpenStudy (anonymous):

we have to agree on what counts as a success.

OpenStudy (anonymous):

No, I dont think so, u will get same answer by more complicated routye....

OpenStudy (anonymous):

hold on. if we do not agree on what a success is then we cannot arrive at the same answer

OpenStudy (anonymous):

1/3 is defining success as rolling odd heads followed by tail.

OpenStudy (anonymous):

believe me i am not saying i am right and you are wrong, i am just saying that we have to know what a success looks like

OpenStudy (anonymous):

If u define success your way, it will not be a third.....

OpenStudy (anonymous):

but i don't think so

OpenStudy (anonymous):

But then u will not multiply by 2 at the end either....

OpenStudy (anonymous):

right. if i count t t h h h t as a success in 6 rolls and you only count h h h h h t, then we cannot come up with same answer

OpenStudy (anonymous):

We will come up with the same answer, the calculations will be different.

OpenStudy (anonymous):

i see 4 ways to get a string of odd number of heads followed by a tail in 6 tosses h h h h h t t t h h h t t t t t h t h h t t h t

OpenStudy (anonymous):

I get 1/3 because tail after even heads is fail in that scenario.....

OpenStudy (anonymous):

really? ok then please go slow because i believe you

OpenStudy (anonymous):

but only in that calculation, it is just a calculation, an abstraction if u will...

OpenStudy (anonymous):

no fine, i just want to see the calculation. i am more or less sure you are right, i just would like to check

OpenStudy (anonymous):

SO probabilty of HT is a half squared right?

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

so maybe we start with x = 1/4 + ...

OpenStudy (anonymous):

and then 3 heads and a tail is to the fourth...

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

what about t h t

OpenStudy (anonymous):

to me that counts and you are ignoring it it seems

OpenStudy (anonymous):

I am only talking about the probability of getting hhht....

OpenStudy (anonymous):

but what about t h t? will we come back to it? because the way i read the problem that counts as a success

OpenStudy (anonymous):

if you jump right to h h h t then i understand exactly what you are going to get and if then next one is h h h h h t etc then i agree with you. but i count other ones

OpenStudy (anonymous):

I am assuming (possibly wrongly) that I will immediately start tossing heads. Let me think about it some more, the trials fail after a tail, the rest is picked up in the 2 at the end.

OpenStudy (anonymous):

Also, u have yet to pick any holes in the algebraic formulation...

OpenStudy (anonymous):

wow i just googled this problem and it is not so straight forward it seems.

OpenStudy (anonymous):

although the answer might in fact be 6 (!)

OpenStudy (anonymous):

I WAS certain about it, now u have made be have doubts..:-) Would like to be certain again....

OpenStudy (anonymous):

my problem is not with the formula. my problem is with ignoring the possibility of t h t being a success. if you are going to compute P(ht) + p(h h h t) + ... you will certainly be right

OpenStudy (anonymous):

My idea is the tail flips are being picked up by the 2, which is also based on three trials ie a 1/3.

OpenStudy (anonymous):

If it isn't a third because some other type of calculation, then the number of trials will be different and so will the 2.

OpenStudy (anonymous):

Going to watch the box and give my head a rest....)

OpenStudy (anonymous):

ok but here is something i nicked from the web. i will read it carefully and see if it makes any sense

OpenStudy (anonymous):

How about the easier formula H=(1/2)*(H+1)+(1/4)*2+(1/4)*(H+2), where H is the expected number of flips? I should not be the one to explain it, since I am borrowing the method from luv2fap, but I will, anyway. Let the expected number of tosses be H. Then, if on the first flip we got T, then we “wasted” one flip, and due to the independence of the tosses, we again expect to stop flipping after H turns (the probability of flipping a T is 1/2). Now, if we managed to flip an H, we need to examine the second toss more carefully. In case of a T – we are done and we used exactly 2 flips (the probability, however, to have HT is 1/4 and this is why we multiply 2 by it). In case of an H – we are back to square 1 having wasted 2 flips and expecting to be done after H flips. The probability for HH is 1/4, again.

OpenStudy (anonymous):

this certainly gives h = 6 for sure. hope i didn't confuse matters too much, but the reasoning is different i think

OpenStudy (anonymous):

How about the easier formula H=(1/2)*(H+1)+(1/4)*2+(1/4)*(H+2) That is the algebraic formula I gave u before...

OpenStudy (anonymous):

yes so it is. sorry for wasting your time. i thought you got this from summing up, but i was mistaken

OpenStudy (anonymous):

No, it was interesting discussion, time not wasted...:-)

OpenStudy (anonymous):

now i that i looked at the explanation above i see where the formula came from. i wonder why this is the same as just adding the geometric series you get for h t h h h t h h h h h t etc and then doubling? good work

OpenStudy (anonymous):

It's not just doubling...:-) (Well, it is but it isn't...)